OpenStudy (anonymous):

n + 5 −16 = −1??? Im confused what am i supposed to do?

OpenStudy (acxbox22):

n+5-16=-1 is that the equation

OpenStudy (anonymous):

no its n + 5 ______ = -1

OpenStudy (anonymous):

n+ 5 _____ = -1 -16

OpenStudy (anonymous):

im dont know what to do

OpenStudy (acxbox22):

start by multiplying by -16 on both sides